# Online computation of normalized substring complexity

### v1

@@ 1 Introduction
Compressibility of a string is a classical way to evaluate its complexity. Compressibility has also other applications such as measuring similarity of two strings, based on the assumption that two strings are similar if the compressibility of one is not much affected by the availability of the other [11]. The ultimate compressibility measure is provided by Kolmogorov complexity which, however, is not practical in any way: it is not computable and is defined up to an additive constant and therefore is not meaningful for an individual finite string. A common practical approach is then to apply conventional string compressors based on efficient compression/decompression algorithms.
@@
In this paper, we study the problem of computing $\delta$ in the online mode. In light of the useful properties of $\delta$, it is important to be able to compute it online, especially for large input strings. Note that online implementation of compression algorithms is an extensively studied issue. In particular, a number of works has been done on efficient computation of Lempel-Ziv parsing, including online computation [27, 33, 36, 29]. Since $\delta$ is a good estimator of the Lempel-Ziv compression measure (cf above), computing it online is an interesting problem which, to our knowledge, has not been studied before. In this paper, we propose two algorithms for computing $\delta$ online: one working in amortized $O(\log n)$ time and another in worst-case $O(\log^{3}n)$ time per character.
@@ 2 Preliminaries
Let $\sigma\geq 2$. For every $w$, $|w|>3$, it holds $\delta\leq n/\log_{\sigma}n$.
@@ 4 Updating the substring count array online
Let us now focus on how the array $c_{w}$ can be modified when a letter $x$ is appended to $w$ on the right. Let $|w|=n$ and $c_{w}[1..n]$ be the substring count array of $w$. We are interested in $c_{wx}[1..n+1])$. The following Lemma specifies the updates.
@@
Appending $x$ to $w$ introduces $n+1$ new substrings which are suffixes of $wx$ of length $1$ to $n+1$. Recall that $\alpha(wx)$ is the shortest of them that has no other occurrences in $wx$. Therefore, the shorter suffixes of length $1$ to $\alpha(wx)-1$ are already present and the counts for $k=1..\alpha(wx)-1$ remain unchanged. For every $k\in[\alpha(wx)..n+1]$, exactly one new substring is introduced.
@@
(i) follows from the definition of $\alpha$. (ii) follows from Theorem 3.1 since $L+1<\alpha_{i+1}\leq R$ would contradict that $c_{w[1..i]}$ must strictly increase on the first interval (property (i) of Theorem 3.1).
Observe that the case $\alpha_{i+1}=R+1$ extends the constant interval ((ii) of Theorem 3.1) by $1$ (i.e. $R$ gets incremented by $1$), the case $\alpha_{i+1}=L+1,L_{i}<R_{i}$ shortens the constant interval by 11 ($L$ gets incremented by $1$), and the remaining cases keep $L$ and $R$ unchanged. The three cases are illustrated in Example 2.
@@
$\sum\{(\alpha_{i-1}-\alpha_{i})\,|\,\alpha_{i}\leq\alpha_{i-1}\}\leq n$.
@@
$\sum\{\alpha_{i}\,|\,\alpha_{i}\leq\alpha_{i-1}\}=O(n\log n)$.
Proof
Consider step $i$. At the previous step i1i-1, $w[i-\alpha_{i-1}..i-1]$ is the shortest suffix that does not occur earlier, and $w[i-\alpha_{i-1}+1..i-1]$ is the longest suffix that does occur earlier. Note that that earlier occurrence must be preceded by another letter than $w[i-\alpha_{i-1}]$, i.e. $w[i-\alpha_{i-1}..i-1]$ is left-special [^1] The earlier occurrence can also be a prefix, therefore we assume that any prefix is preceded by a special letter that does not occur in the rest of $w$. In other words, we assume that $w$ starts with a special unique letter.. When letter $w[i]$ is processed, a pullback occurs when $\alpha_{i}\leq\alpha_{i-1}$ holds, that is, when the shortest unique suffix is strictly shorter than $\alpha_{i-1}+1$. This implies that $w[i-\alpha_{i-1}..i]$ does not occur earlier, that is any earlier occurrence of $w[i-\alpha_{i-1}..i-1]$ is followed by another letter than $w[i]$. This shows that $w[i-\alpha_{i-1}..i-1]$ is either a new bispecial substring or an existing bispecial substring with a new right context. Observe that this new occurrence of a bispecial substring is permanent, that is, it will persist when new letters are appended to $w$.
In[16] the authors study irreducible LCP (longest common prefix) values and prove that their sum is bounded by $2n\log n$. An LCP value is the length of the longest common prefix between a suffix and the lexicographically preceding one.An LCP value is irreducible if these two suffixes are preceded by distinct letters. This implies that each irreducible LCP value corresponds to a pair of occurrences of a bispecial substring, except if one of the involved suffixes if a prefix of the other one. Inversely, each newly appeared occurrence of a bispecial substring corresponds to an irreducible LCP. If the same bispecial substring $u$ has $q$ occurrences, it will contribute with q1q-1 irreducible LCP values $|u|$.
We conclude that when a pullback occurs at step $i$, we get a new irreducible LCP value of $\alpha_{i-1}$. From [16, Theorem 1], we obtain the result.
We remark that the sum of irreducible LCP values can be bounded more precisely in terms of compression measures: thus, it is proven in [17] that this sum is $O(n\log\delta)$. This bound holds for the sum of $\alpha_{i}$ as well.
@@ 5.2 Convex hull reduction
The ratio $c[k]/k$ is the slope of the line between the origin and $(k,c[k])$. Therefore, $\delta(w)$ corresponds to the point $(k^{\prime},c[k^{\prime}])\in\mathcal{P}$ with the maximal slope. Clearly, $(k^{\prime},c[k^{\prime}])$ belongs to the upper convex hall and, on the other hand, to the tangent line from the origin to this convex hull.
@@ 5.3 Maintaining a point of maximum slope
The algorithm implied by update (Sect. 4) is a sequence of $n$ steps, so that at each step $i$ we (i) either increment the values in some contiguous right part $c[\alpha_{i}..R_{i}]$ of the current array $c$, or (ii) extend array $c$ on the right by duplicating the last entry. Along this process, we have to maintain $\max_{1\leq k\leq R_{i}}\{c[k]/k\}$. We call the active interval the fragment $c[\alpha_{i}..R_{i}]$ that is incremented at the current step $i$. Note that for operation (ii), the active interval is $c[R_{i}..R_{i}]$, as this case occurs when $\alpha_{i}=R_{i-1}+1$ (see Section 4). For an interval $\mathcal{I}\subseteq[1..R_{i}]$, we say that $\widetilde{k}_{\mathcal{I}}=\operatorname*{arg\,max}_{k\in\mathcal{I}}\{c[k]/k\}$ is the leader on the interval $\mathcal{I}$, or simply the leader if $\mathcal{I}=[1..R_{i}]$.
@@
We dynamically maintain the active interval $c[\alpha_{i}..R_{i}]$ and the interval $c[1..\alpha_{i}-1]$ before the active one, in two different ways. For the interval $c[1..\alpha_{i}-1]$, we will explicitly store values $\frac{c[k]}{k}$ and maintain the leader $\operatorname*{arg\,max}_{1\leq k\leq\alpha_{i}-1}\frac{c[k]}{k}$. To this aim, we use two arrays $\mathit{val}$ and $\mathit{maxval}$ where $\mathit{val}[k]=\frac{c[k]}{k}$ and $\mathit{maxval}[k]=\max_{1\leq j\leq k}\frac{c[j]}{j}$. Since within each round, we have $\alpha_{i}=\alpha_{i-1}+1$, upating $\mathit{val}$ and $\mathit{maxval}$ takes $O(1)$ operations at each step. For the pullback step, no updates of $\mathit{val}$ and $\mathit{maxval}$ are needed.
For the active interval $c[\alpha_{i}..R_{i}]$, we maintain the leader using the reduction to the convex hull problem explained earlier. We will maintain the convex hull of the active interval using the data structure of Theorem 5.2. However, each point $(k,c[k])$ of the active interval will be represented by point $(k,c[k]-\Delta)$, for a dynamic parameter $\Delta$ called the active shift. The meaning of $\Delta$ is the vertical shift of the actual point $(k,c[k])$ relative to the $y$-coordinate of the point currently stored in the data structure. Our algorithm will insure that the value of $\Delta$ is the same for all points in the active interval of the array.
@@
For any $k$ in the interval $1..\alpha_{i}-1]$, all values $\mathit{val}[k]$ and $\mathit{maxval}[k]$ are correct.
At initialization, after the first character of the text is read, we have $c[1]=1$, $\alpha_{1}=R_{1}=1$, and we set $\mathit{val}[1]=1$, $\mathit{maxval}[1]=1$, $\Delta=0$, and insert point $(1,c[1])$ into the convex hull structure.
At each step $i\geq 2$, we proceed as follows. First observe that operation (ii) is easily implemented by retrieving the rightmost point $(R_{i-1},c[R_{i-1}])$ from the convex hull data structure [^4] Retrieval of extreme points is trivial as we can simply explicitly maintain them in separate variables and inserting a new rightmost point $(R_{i},c[R_{i-1}])$ for $R_{i}=R_{i-1}+1$. Consider now operation (i). If $\alpha_{i}=\alpha_{i-1}+1$, we retrieve the current leftmost point $(\alpha_{i-1},c[\alpha_{i-1}]-\Delta)$ from the convex hull data structure and determine the value $c[\alpha_{i-1}]$, update $\mathit{val}$ and $\mathit{maxval}$ by $\mathit{val}[\alpha_{i-1}]\leftarrow c[\alpha_{i-1}]/\alpha_{i-1}$ and $\mathit{maxval}[\alpha_{i-1}]\leftarrow\max(\mathit{val}[\alpha_{i-1}],\mathit{maxval}[\alpha_{i-1}-1])$, then remove the leftmost point $(\alpha_{i-1},c[\alpha_{i-1}]-\Delta)$ from the convex hull data structure and increment $\Delta$ by $1$. Otherwise, if $\alpha_{i}\leq\alpha_{i-1}$, we execute the pullback step. In this case, we first increment $\Delta$ by $1$ and then consecutively insert points $(k,c[k]-\Delta)$ for all $k$ from $\alpha_{i-1}-1$ down to $\alpha_{i}$ into the convex hull data structure. We observe that every inserted point $(k,c[k]-\Delta)$ is to the left of all points already stored in the data structure.
@@
The Lemma follows by induction on steps. Invariant 1 is verified at initialization. Suppose it is satisfied after (i1)(i-1) steps for some $i>1$. If step $i$ is a forward step, the claim is straightforward. If step $i$ is a pullback step, we insert points $(k,c[k]-\Delta)$ for all $k$ such that $\alpha_{i}\leq k<\alpha_{i-1}$, which guarantees that the active shift remains the same for all points in the data structure.
@@
At each step, we can compute the leader by taking the maximum of leader values on the two intervals. Given a set $P$ of planar points and $d>0$, we will say that the set $P^{\prime}$ is a vertical $d$-shift of $P$ iff for every point $p=(x,y)\in P$ there is exactly one point $p^{\prime}=(x-d,y)\in P^{\prime}$. To compute the leader on the active interval, we answer a tangent query to the convex hull, based on the following straightforward observation.
@@
Let $\mathcal{P}_{a}=\{(k,c[k]),k=\alpha_{i}..R\}$ be the set of points of the current active interval. By Lemma 6, the set of points $\mathcal{P}^{\prime}_{a}$, stored in our convex hull structure, is a vertical $\Delta$-shift of $\mathcal{P}$. Let $p^{\prime}_{a}$ be the tangent point of $(0,-\Delta)$ and the convex hull of $\mathcal{P}^{\prime}_{a}$. Let $p_{a}$ be the tangency point from $(0,0)$ to the convex hull of $\mathcal{P}_{a}$. By Lemma 8 $p$ and $p_{a}$ have the same $x$-coordinates and their $y$-coordinates differ by $\Delta$: if $p^{\prime}_{a}=(\widetilde{k}_{a},\nu)$, then $p_{a}=(\widetilde{k}_{a},\nu+\Delta)$. Hence $\widetilde{k}_{a}$ is the leader in the active interval and its value is $\nu+\Delta$. By Theorem 5.2, we can find $p^{\prime}_{a}$ in $O(\log n)$ time using our convex hull data structure. The resulting leader value after step $i$ is equal to $\max(\mathit{maxval}[\alpha_{i}-1],(\nu+\Delta)/\widetilde{k}_{a})$.
@@ 5.4 Worst-case solution
Now we explain how we modify the data structure of  [28] to support our operations (i) and (ii). The main difference is that we do not store in $\mathcal{T}$ the $y$-coordinates of points: every point $(k,c[k])$ is represented in $\mathcal{T}$ by its $x$-coordinate $k$ only. $y$-coordinates $c[k]$ are stored in a separate dynamic range tree $\mathcal{T}_{c}$. For every point $(k,c[k])$, there is a leaf of $\mathcal{T}_{c}$ that holds it. Similar to $\mathcal{T}$, leaves of $\mathcal{T}_{c}$ are sorted by $x$-coordinates of points. Each time we need to retrieve the $y$-coordinate of some point $(k,c[k])$ we use $\mathcal{T}_{c}$ for this purpose.

### v2

@@ 1 Introduction
Compressibility of a string is a classical way to evaluate its complexity. Compressibility has also other applications such as measuring similarity of two strings, where the similarity is expressed by how well one of the strings can be compressed if the other is provided as a reference [11]. The ultimate compressibility measure is provided by Kolmogorov complexity which, however, is not practical in any way: it is not computable and is defined up to an additive constant and therefore is not meaningful for an individual finite string. A common practical approach is then to apply conventional string compressors based on efficient compression/decompression algorithms.
@@
In this paper, we study the problem of computing $\delta$ in the online mode. In light of the useful properties of $\delta$, it is important to be able to compute it online, especially for large input strings. Note that online implementation of compression algorithms is an extensively studied issue. In particular, a number of works has been done on efficient computation of Lempel-Ziv parsing, including online computation [27, 33, 36, 29]. Since $\delta$ is a good estimator of the Lempel-Ziv compression measure (cf above), maintaining its exact value of online is an interesting problem which, to our knowledge, has not been studied before. A related work includes [2] where a probabilistic sketch-based streaming algorithm for computing an approximation of $\delta$ has been studied. Another related work is [32] which focuses on a probabilistic sublinear (non-online) algorithm for approximating $\delta$. In this paper, we propose two algorithms for computing the exact value of $\delta$ online: one working in amortized $O(\log n)$ time and another in worst-case $O(\log^{3}n)$ time per character.
@@ 2 Preliminaries
Let $\sigma\geq 2$. For every $w$, $n>3$, it holds $\delta\leq n/\log_{\sigma}n$.
@@ 4 Updating the substring count array online
Let us now focus on how the array $c_{w}$ can be modified when a letter $x$ is appended to $w$ on the right. Let $c_{w}[1..n]$ be the substring count array of $w$. We are interested in $c_{wx}[1..n+1]$. The following Lemma specifies the updates.
@@
Appending $x$ to $w$ introduces $n+1$ potentially new substrings which are suffixes of $wx$ of length $1$ to $n+1$. By definition, $\alpha(wx)$ is the length of the shortest suffix that has no other occurrences in $wx$. Therefore, the shorter suffixes of length $1$ to $\alpha(wx)-1$ are already present and the counts for $k=1..\alpha(wx)-1$ remain unchanged. For every $k\in[\alpha(wx)..n+1]$, exactly one new substring is introduced.
@@
(i) follows from the definition of $\alpha$, as the shortest non-repeated suffix of $w[1..i+1]$ cannot be more than by $1$ longer than that of $w[1..i]$. To show (ii), first recall that by Theorem 3.1(v),(iv), we have either $\alpha_{i}\leq L$ or $\alpha_{i}=R$ (or both, if $L=R$). If $\alpha_{i}\leq L$, then $\alpha_{i+1}\leq L+1$ follows from (i). Consider the case $\alpha_{i}=R$. $L+1<\alpha_{i+1}\leq R$ would contradict the structure of the array $c_{w[1..i+1]}$ implied by Theorem 3.1, namely that $c_{w[1..i+1]}$ must strictly increase on the first interval (Theorem 3.1(i)). Therefore, we must have $\alpha_{i+1}=R+1$ or $\alpha_{i+1}\leq L+1$.
Observe that the case $\alpha_{i+1}=R+1$ extends the constant interval ((ii) of Theorem 3.1) by $1$ (i.e. $R$ gets incremented by $1$), the case $\alpha_{i+1}=L+1,L<R$ shortens the constant interval by 11 ($L$ gets incremented by $1$), and the remaining cases keep $L$ and $R$ unchanged. The three cases are illustrated in Example 2.
@@
$\sum_{i\,|\,\alpha_{i}\leq\alpha_{i-1}}(\alpha_{i-1}-\alpha_{i})\leq n$.
@@
$\sum_{i\,|\,\alpha_{i}\leq\alpha_{i-1}}\alpha_{i}=O(n\log n)$.
The proof of Theorem 4.1 is given in the Appendix.
@@ 5.2 Convex hull reduction
The ratio $c[k]/k$ is the slope of the line between the origin and $(k,c[k])$. Therefore, $\delta(w)$ corresponds to the point $(k^{\prime},c[k^{\prime}])\in\mathcal{P}$ with the maximal slope. Clearly, $(k^{\prime},c[k^{\prime}])$ belongs to the upper convex hull and, on the other hand, to the tangent line from the origin to this convex hull.
@@ 5.3 Maintaining a point of maximum slope
The algorithm implied by update (Sect. 4) is a sequence of $n$ steps, so that at each step $i$ we (i) either increment the values in some contiguous right part $c[\alpha_{i}..R_{i}]$ [^3] In this section, we use notation $R_{i}$ to refer to the value of $R$ at step $i$ of the current array $c$, or (ii) extend array $c$ on the right by duplicating the last entry. Along this process, we have to maintain $\max_{1\leq k\leq R_{i}}\{c[k]/k\}$. We call the active interval the fragment $c[\alpha_{i}..R_{i}]$ that is incremented at the current step $i$. Note that for operation (ii), the active interval is $c[R_{i}..R_{i}]$, as this case occurs when $\alpha_{i}=R_{i-1}+1$ (see Section 4).
For an interval $\mathcal{I}\subseteq[1..R_{i}]$, we say that $\widetilde{k}_{\mathcal{I}}=\operatorname*{arg\,max}_{k\in\mathcal{I}}\{c[k]/k\}$ is the leader on the interval $\mathcal{I}$, or simply the leader if $\mathcal{I}=[1..R_{i}]$.
@@
We dynamically maintain the active interval $c[\alpha_{i}..R_{i}]$ and the interval $c[1..\alpha_{i}-1]$ before the active one, in two different ways. For the interval $c[1..\alpha_{i}-1]$, we will explicitly store values $c[k]$ and maintain the leader $\operatorname*{arg\,max}_{1\leq k\leq\alpha_{i}-1}\frac{c[k]}{k}$. To this aim, we use two arrays $\mathit{val}$ and $\mathit{leader}$ where $\mathit{val}[k]=c[k]$ and $\mathit{leader}[k]=\operatorname*{arg\,max}_{1\leq j\leq k}\frac{val[j]}{j}$. Arrays $\mathit{val}$ and $\mathit{leader}$ are dynamic and may grow in time, we implement them using the standard doubling technique. Since within each round, we have $\alpha_{i}=\alpha_{i-1}+1$, updating $\mathit{val}$ and $\mathit{leader}$ takes $O(1)$ operations at each step. For the pullback step, no updates of $\mathit{val}$ and $\mathit{leader}$ are needed.
For the active interval $c[\alpha_{i}..R_{i}]$, we maintain the leader using the reduction to the convex hull problem explained earlier. We will maintain the convex hull of the active interval using the data structure of Theorem 5.2. However, each point $(k,c[k])$ of the active interval will be represented by point $(k,c[k]-\Delta)$, for a dynamic parameter $\Delta$ called the active shift. The meaning of $\Delta$ is the vertical shift of the actual point $(k,c[k])$ relative to the $y$-coordinate of the point currently stored in the data structure. Our algorithm will ensure that the value of $\Delta$ is the same for all points in the active interval of the array.
@@
For any $k$ in the interval $1..\alpha_{i}-1]$, all values $\mathit{val}[k]$ and $\mathit{leader}[k]$ are correct.
At initialization, after the first character of the text is read, we have $c[1]=1$, $\alpha_{1}=R_{1}=1$, and we insert point $(1,c[1])$ into the convex hull structure, set $\Delta=0$, and initialize empty arrays $\mathit{val}$ and $\mathit{leader}$.
At each step $i\geq 2$, we proceed as follows. First observe that operation (ii) is easily implemented by retrieving the rightmost point $(R_{i-1},c[R_{i-1}]-\Delta)$ from the convex hull data structure [^4] The convex hull data structure described in [6, Theorem 5] supports the retrieval of extreme points in $O(1)$ time. Alternatively, we can maintain a separate list of all points of the active interval sorted by $x$-coordinates. and inserting a new rightmost point $(R_{i},c[R_{i-1}]-\Delta)$ for $R_{i}=R_{i-1}+1$. Consider now operation (i). If $\alpha_{i}=\alpha_{i-1}+1$, we retrieve the current leftmost point $(\alpha_{i-1},c[\alpha_{i-1}]-\Delta)$ from the convex hull data structure and determine the value $c[\alpha_{i-1}]$ by adding $\Delta$, then update $\mathit{val}$ and $\mathit{leader}$. Specifically, we set $\mathit{val}[\alpha_{i-1}]\leftarrow c[\alpha_{i-1}]$ and set $\mathit{leader}[\alpha_{i-1}]\leftarrow\alpha_{i-1}$ if $\mathit{val}[\alpha_{i-1}]/\alpha_{i-1}\geq\mathit{val}[\mathit{leader}[\alpha_{i-1}-1]]/\mathit{leader}[\alpha_{i-1}-1]$ and $\mathit{leader}[\alpha_{i-1}]\leftarrow\mathit{leader}[\alpha_{i-1}-1]$ otherwise. Then we remove the leftmost point $(\alpha_{i-1},c[\alpha_{i-1}]-\Delta)$ from the convex hull data structure and increment $\Delta$ by $1$. Otherwise, if $\alpha_{i}\leq\alpha_{i-1}$, we execute the pullback step. In this case, we consecutively insert points $(k,\mathit{val}[k]-\Delta)$ for all $k$ from $\alpha_{i-1}-1$ down to $\alpha_{i}$ into the convex hull data structure. We observe that every inserted point is to the left of all points already stored in the data structure. Then we increment $\Delta$ by $1$.
@@
The Lemma follows by induction on steps. Invariant 1 is verified at initialization. Suppose it is satisfied after (i1)(i-1) steps for some $i>1$. If step $i$ is an increment step, the claim is straightforward. If step $i$ is a pullback step, we insert points $(k,c[k]-\Delta)$ for all $k$ such that $\alpha_{i}\leq k<\alpha_{i-1}$, which guarantees that the active shift remains the same for all points in the data structure.
@@
At each step, we can compute the leader by taking the maximum of leader values on the two intervals. Given a set $P$ of planar points and $d>0$, we will say that the set $P^{\prime}$ is a vertical $d$-shift of $P$ iff for every point $p=(x,y)\in P$ there is exactly one point $p^{\prime}=(x,y-d)\in P^{\prime}$. To compute the leader on the active interval, we answer a tangent query to the convex hull, based on the following straightforward observation.
@@
Let $\mathcal{P}_{a}=\{(k,c[k]),k=\alpha_{i}..R\}$ be the set of points of the current active interval. By Lemma 6, the set of points $\mathcal{P}^{\prime}_{a}$, stored in our convex hull structure, is a vertical $\Delta$-shift of $\mathcal{P}$. Let $p^{\prime}_{a}$ be the tangent point of $(0,-\Delta)$ and the convex hull of $\mathcal{P}^{\prime}_{a}$. Let $p_{a}$ be the tangency point from $(0,0)$ to the convex hull of $\mathcal{P}_{a}$. By Lemma 8 $p$ and $p_{a}$ have the same $x$-coordinates and their $y$-coordinates differ by $\Delta$: if $p^{\prime}_{a}=(\widetilde{k}_{a},\nu)$, then $p_{a}=(\widetilde{k}_{a},\nu+\Delta)$. Hence $\widetilde{k}_{a}$ is the leader in the active interval and its value is $\nu+\Delta$. By Theorem 5.2, we can find $p^{\prime}_{a}$ in $O(\log n)$ time using our convex hull data structure. The resulting leader after step $i$ is either $\widetilde{k}_{a}$ or $\mathit{leader}[\alpha_{i}-1]$, depending on which of the fractions $(\nu+\Delta)/\widetilde{k}_{a}$ or $\mathit{val}[\mathit{leader}[\alpha_{i}-1]]/\mathit{leader}[\alpha_{i}-1]$ is larger.
@@ 5.4 Worst-case solution
Now we explain how we modify the data structure of  [28] to support our operations (i) and (ii). The main difference is that we do not store in $\mathcal{T}$ the $y$-coordinates of points: every point $(k,c[k])$ is represented in $\mathcal{T}$ by its $x$-coordinate $k$ only. $y$-coordinates $c[k]$ are stored in a separate dynamic range tree $\mathcal{T}_{c}$. For every point $(k,c[k])$, there is a leaf of $\mathcal{T}_{c}$ that corresponds to it. Similar to $\mathcal{T}$, leaves of $\mathcal{T}_{c}$ are sorted by $x$-coordinates of points. Each time we need to retrieve the $y$-coordinate of some point $(k,c[k])$ we use $\mathcal{T}_{c}$ for this purpose.